How many ways can a 7-member board of trustees be created from a group of 10
men and 8 women if 4 members must be male and 3 members must be female?

Answer by math_helper(2461)



# ways to choose the 4 men from 10
times
# ways to choose the 3 women from 8

10C4 * 8C3 = (10*9*8*7)/(4*3*2*1) * (8*7*6)/(3*2*1) = 11760 ways